# Prefix Sums via Kronecker Products

### v1

@@ I Introduction
A plethora of sophisticated prefix circuit constructions have been proposed in the literature with logarithmic depth [1, 9, 2, 10, 11, 12, 13, 14, 15, 16, 17] (see Table 1). Sklansky [1] described a circuit that achieves the optimal depth $D(n)=\log(n)$, but the size increases to $n\log(n)/2$. Ladner and Fischer [9] proposed a recursive construction that can reduce the size by slightly increasing the depth. Brent and Kung [2] described a recursive circuit in the context of parallel adders. Snir [10] proved a tight optimality condition for the sum of depth and size of any prefix circuit, specifically, the following inequality:
@@
where $F(t)$ is the $t$-th Fibonacci number. In addition, they described a circuit family that achieves this lower bound for every $n$ (see also Table 1).
@@
$n\lceil\log(n)\rceil-n+1$ $\log(n)+O(1)$ 22 $O(n\log(n))$ Kogge–Stone [21],[3, Sec. 3.1, Fig. 1].(2)
2n22n-2 $\log(n)+O(1) 22$ $O(\log(n))$ Cyclic Reduction [3, Sec. 3.1, Fig. 5].(2)
@@
2n2D(n)2n-2-D(n) $2\log(n)+O(1)$ $2\lceil\log(n)\rceil-2$ 0 LYD [11], $n\geq 9$.
2n2D(n)2n-2-D(n) $2\log(n)+O(1)$ $\lceil\log(n)\rceil+1$ 0 SL [12], $n\geq 12$.
@@
2n2D(n)2n-2-D(n) $\mathbf{1.4401}\log(n)+O(1)$ $D(n)$ 0 ZCG [18].
2n2D(n)2n-2-D(n) $2\log(n)+O(1)$ $\mathbf{3}$ 0 This work: Thm. 2 for s=2s=2.(2)
2n2D(n)2n-2-D(n) $\mathbf{1.8928}\log(n)+O(1)$ 44 0 This work: Thm. 2 for s=3s=3.(2)
@@
By carefully organizing the underlying operations, we obtain a new family of recursive prefix circuits, whose properties are summarized in Table 1 (see also Theorem 2). These circuits are parametrized by an integer $s\in[2,n/2]$ which corresponds to the “block-size” of the recursion. Different choices for $s$ yield circuits with different properties. We highlight the following cases:
  • For $s=2$, the proposed circuit has the lowest known fan-out compared to all known zero-deficiency constructions with depth $2\log(n)+O(1)$.
  • For $s=3$, it has the lowest depth $(\approx 1.893\log(n)+O(1))$ between all circuits with constant fan-out.
Importantly, the proposed circuits are quite simple to describe. Specifically, they satisfy the so-called LOGTIME-uniformity, which means that the connections between the nodes of a circuit can be verified in $O(\log(n))$ time for a sequence of length $n$ [22]. To the best of our knowledge this is the first LOGTIME-uniform prefix circuit family that achieves depth smaller than $2\log(n)+O(1)$.
@@ II Kronecker products
We frequently use the following two Kronecker properties, see [23, Equations $2.4$ and $2.10$]:
@@ 3. Step 3:
For every pair of integers $n\geq 4$ and $s\in[2,n/2]$, there exists a LOGTIME-uniform family of recursive zero-deficiency prefix circuits with depth $sh_{s}^{*}(n)+r_{s}(n)-1\leq s\lceil\log_{s}(n)\rceil-2$ and fan-out $s$. If $n$ is a power of $s$, there is a single gate with fan-out $s+1$.
@@
We show by induction that a recursive construction achieves zero-deficiency. In the base case, if $\lceil\tfrac{n}{s}\rceil-1\leq s$, then we simply have a single serial prefix circuit with size n1n-1 and depth n1n-1, which has zero-deficiency. For the inductive step, we argue that if the claim holds for length $\lceil\tfrac{n}{s}\rceil-1$, then it also holds for length $n$. The first layer has size $n-\lceil\tfrac{n}{s}\rceil$ and depth s1s-1. The circuit of the second layer has zero-deficiency by the induction hypothesis, and therefore its depth and size sum to $2(\lceil\tfrac{n}{s}\rceil-1)-2=2\lceil\tfrac{n}{s}\rceil-4$. The final layer has depth one and size $n-s-\lceil\tfrac{n}{s}\rceil+2$. Adding everything together, the total depth and size sum to:
$\overbrace{n-\lceil\tfrac{n}{s}\rceil+s-1}^{\text{Layer 1}}+\overbrace{2\lceil\tfrac{n}{s}\rceil-4}^{\text{Layer 2}}+\overbrace{n-s-\lceil\tfrac{n}{s}\rceil+3}^{\text{Layer 3}}=2n-2,$
as desired. We can now upper bound the depth of the recursive circuit. We have that:
@@
This quantity is always upper bounded by $s\lceil\log_{s}(n)\rceil-2$. To see this, note that $D(n)$ is equal to $sh^{*}_{s}(n)+r_{s}(n)-1$. Now, it always holds that $D(n)\leq D(s^{k})$, where $k=\arg\min\{j\ |\ s^{k}\geq n\}$. But if we replace $n$ with $s^{k}$ in the above, it holds that $h^{*}_{s}(s^{k})=k-1$ and $r_{s}(s^{k})=h_{s}^{(k-1)}=s-1$. Therefore, the total depth satisfies $D(s^{k})=s(k-1)+(s-1)-1=sk-2$. Since $k=\lceil\log_{s}(n)\rceil$, we finally obtain that $D(n)\leq D(s^{k})=s\lceil\log_{s}(n)\rceil-2,$ where the equality always holds when $n$ is a power of $s$.
To see that the construction can be verified in LOGTIME, note that given $n$, $s$, $i\in[n]$, $j\in[n]$, and $k\in[s\lceil\log_{s}(n)]$, as inputs, it is straightforward to answer in $O(\log(n))$ steps whether input $i$ is connected with input $j$ at level $k$. ∎
@@
For $s=2$, we obtain a zero-deficiency circuit with depth at most $2\lceil\log(n)\rceil-2\in 2\log(n)+O(1)$, which is the same as Brent–Kung for powers of two. However, $s=2$ does not always give the minimum achievable depth. Indeed, for larger values of $n$ we can sharpen the bound by setting $s=3$, which gives $D(n)\leq 3\lceil\log_{3}(n)\rceil-2\in\frac{3}{\log(3)}\log(n)+O(1)\approx 1.8928\log(n)+O(1)$. For any $n$, the actual minimum depth can be computed with dynamic programming: $\min D(n)=\min_{2\leq s\leq n/2}\{s+\min D(\lceil\frac{n}{s}\rceil-1)\}$.
@@ IV Quantum adders
Here we mention how to apply the proposed prefix circuits to quantum adders. Such circuits have received significant attention the last three decades since they serve as building blocks for important problems such as discrete logarithm and integer factoring [27, 28, 29, 30, 31, 32, 4, 5]. Here we follow the methodology of [4, 5], where the main idea is to use a prefix circuit for carry propagation/generation (carry look-ahead framework). Circuits are evaluated based on the following metrics:
@@
The maximum number of Toffoli gates in any path from an input to an output.
@@
The goal is to minimize the Toffoli depth/count by carefully overlapping different independent Toffoli layers, while still using as few auxiliary qubits as possible. Figure 3 shows three cases where Toffoli gates are used in the quantum circuit for propagation and generation. We refer also to [5, Sections 3 and 4] for a broad overview of the techniques and details of the specific operations.
@@ 3. $k:\bm{g}(i)\leftarrow\mathcal{T}\Big(\bm{g}(i-1),\bm{p}_{t-1}(i),\bm{g}(i)\Big)$ $i\in[s-1:s:n-1]$
The total number of Toffoli gates is bounded as follows. In the first two propagation layers, $\bm{p}_{0}$ and $\bm{p}_{1}$, there are a total of $3n/2$ gates, and additional $n/2$ gates for the un-computation of $\bm{p}_{1}$. In each recursive step $t=1,\ldots,\lceil\log_{s}(n)\rceil-1$, there are at most $n/s^{t}$ gates for $\bm{g}^{(t+1)}$, at most $n/s^{t+1}$ gates for $\bm{p}^{(t+2)}$, at most $n/2^{t}$ gates for $\bm{g}^{(t)}$, and finally at most $n/2^{t+1}$ gates for the uncomputation of the $\bm{p}^{(t+2)}$. This gives a total of at most $3n/2^{t}$ gates at level $t$. Summing for all $t=1,\ldots,\lceil\log(n)\rceil-1$, we obtain at most $3n$ Toffoli gates. ∎

### v2

@@ I Introduction
The simplest way to solve the prefix problem is with a serial circuit, which updates each output as $\bm{y}(i)\leftarrow\bm{x}(i)\circ\bm{y}(i-1)$, with size n1n-1 and depth n1n-1. A plethora of sophisticated prefix circuits have been proposed in the literature with logarithmic depth (see also Table 1). Sklansky [25] described a circuit that achieves the optimal depth $D(n)=\log(n)$, but the size increases to $n\log(n)/2$. Ladner and Fischer [13] proposed a recursive construction that can reduce the size by slightly increasing the depth. Fich [8] proposed improvements and also proved lower bounds for the size of circuits of minimal depth. Brent and Kung [4] described a recursive circuit in the context of parallel adders. Snir [26] proved a tight optimality condition for the sum of depth $D(n)$ and size $S(n)$ of any prefix circuit:
@@
where $F(t)$ is the $t$-th Fibonacci number. In addition, they described a circuit family that achieves this lower bound for every $n$. Sheeran and Parberry [27] proposed a circuit family with parametrized fan-out, which yields a circuit with depth $2\log(n)-1$, zero-deficiency, and fan-out two. Lin and Hung [15] proposed an alternative construction with the same properties. Sergeev [23] showed that the $2\log(n)+O(1)$ depth of [27, 15] is essentially optimal for zero-deficiency circuits with fan-out two, and also proposed new constructions that achieve it.
@@
$n\lceil\log(n)\rceil-n+1$ $\log(n)+O(1)$ 22 $O(n\log(n))$ Kogge–Stone [10],[12, Sec. 3.1, Fig. 1].(2)
2n22n-2 $\log(n)+O(1) 22$ $O(\log(n))$ Cyclic Reduction [12, Sec. 3.1, Fig. 5].(2)
@@
2n2D(n)2n-2-D(n) $2\log(n)+O(1)$ $2\lceil\log(n)\rceil-2$ 0 Lakshmivarahan–Yang–Dhall [20], $n\geq 9$.
2n2D(n)2n-2-D(n) $2\log(n)+O(1)$ $\lceil\log(n)\rceil+1$ 0 Lin–Shih [18], $n\geq 12$.
@@
2n2D(n)2n-2-D(n) $\mathbf{1.4401}\log(n)+O(1)$ $D(n)+1$ 0 Zhu–Cheng–Graham [34].
2n2D(n)2n-2-D(n) $\log(n)+O(1) 𝟐\mathbf{2}$ 0 Sheeran–Parberry [27], Lin–Hung [15], Sergeev [23].
2n2D(n)2n-2-D(n) $\log(n)+O(1) 𝟐\mathbf{2}$ 0 This work: Thm. 2 for s=2s=2.(2)
2n2D(n)2n-2-D(n) $\mathbf{1.8928}\log(n)+O(1)$ 33 0 This work: Thm. 2 for s=3s=3.(2)
@@
By carefully organizing the underlying operations, we obtain a new family of recursive prefix circuits, whose properties are summarized in Table 1 (see also Theorem 2). These circuits are parametrized by an integer $s\in[2,n/2]$ that corresponds to the “block-size” of the recursion. We highlight the following:
  • For $s=3$, it has the lowest depth $(\approx 1.893\log(n)+O(1))$ among all zero-deficiency circuits with constant fan-out.
Importantly, the proposed circuits satisfy the so-called LOGTIME-uniformity, which means that the connections between the nodes of a circuit can be verified in $O(\log(n))$ time for a sequence of length $n$ [28]. In that sense, it is the simplest zero-deficiency circuit with depth $2\log(n)+O(1)$ and fan-out two ($s=2$), and, to the best of our knowledge, it is the first LOGTIME-uniform zero-deficiency prefix circuit family with constant fan-out that achieves depth smaller than $2\log(n)+O(1)$ $(s=3)$.
@@ II Kronecker products
We frequently use the following two Kronecker properties, see [21, Equations $2.4$ and $2.10$]:
@@ 3. Step 3:
For every pair of integers $n\geq 4$ and $s\in[2,n/2]$, there exists a LOGTIME-uniform family of recursive zero-deficiency prefix circuits with depth at most $sh_{s}^{*}(n)+r_{s}(n)\leq s\lceil\log_{s}(n)\rceil-1$ and fan-out $s$.
@@
We first show by induction that a recursive construction achieves zero-deficiency. In the base case, if $\lceil\tfrac{n}{s}\rceil-1\leq s$, then we simply have a single serial prefix circuit with size n1n-1 and depth n1n-1, which has zero-deficiency. For the inductive step, we argue that if the claim holds for length $\lceil\tfrac{n}{s}\rceil-1$, then it also holds for length $n$. The first layer has size $n-\lceil\tfrac{n}{s}\rceil$ and depth s1s-1. The circuit of the second layer has zero-deficiency by the induction hypothesis, and therefore its depth and size sum to $2(\lceil\tfrac{n}{s}\rceil-1)-2=2\lceil\tfrac{n}{s}\rceil-4$. The final layer has depth one and size $n-s-\lceil\tfrac{n}{s}\rceil+2$. Adding everything together, the total depth and size sum to the desired 2n22n-2:
$\overbrace{n-\lceil\tfrac{n}{s}\rceil+s-1}^{\text{Layer 1}}+\overbrace{2\lceil\tfrac{n}{s}\rceil-4}^{\text{Layer 2}}+\overbrace{n-s-\lceil\tfrac{n}{s}\rceil+3}^{\text{Layer 3}}=2n-2.$
We can now upper bound the depth of the recursive circuit. First, we assume that $n$ is not a power of $s$, which is addressed later. Note that after the first recursion, the number of inputs in each recursive step is never a power of $s$, even if the original $n$ is. We have that:
@@
This quantity is always upper bounded by $s\lceil\log_{s}(n)\rceil-2$. To see this, note that $D(n)$ is equal to $sh^{*}_{s}(n)+r_{s}(n)-1$. Now, it always holds that $D(n)\leq D(s^{k})$, where $k=\arg\min\{j\ |\ s^{k}\geq n\}$. But if we replace $n$ with $s^{k}$ in the above, it holds that $h^{*}_{s}(s^{k})=k-1$ and $r_{s}(s^{k})=h_{s}^{(k-1)}=s-1$. Therefore, the total depth satisfies $D(s^{k})=s(k-1)+(s-1)-1=sk-2$. Since $k=\lceil\log_{s}(n)\rceil$, we finally obtain that $D(n)\leq D(s^{k})=s\lceil\log_{s}(n)\rceil-2$.
So far, the maximum fan-out of any node is at most $s$. However, when $n$ is a power of $s$, there is always a single gate at the last level that has fan-out $s+1$. This can be reduced to $s$ simply by constructing a circuit for the first n1n-1 inputs (where n1n-1 is no longer a power of $s$) and then attaching a single gate that adds the $n$-th input to the (n1)(n-1)-th output. This increases the depth by $1$, and reduces the maximum fan-out from $s+1$ to $s$. The circuit retains zero-deficiency.
To see that the construction can be verified in LOGTIME, note that given $n$, $s$, $i\in[n]$, $j\in[n]$, and $k\in[s\lceil\log_{s}(n)-1]$, as inputs, it is straightforward to answer in $O(\log(n))$ steps whether there is a connection between node $i$ at level $k$ and node $j$ at level $k+1$. ∎
@@
For $s=2$, we obtain a zero-deficiency circuit with depth at most $2\lceil\log(n)\rceil-1\in 2\log(n)+O(1)$, which is the same as Brent–Kung for powers of two. However, $s=2$ does not always give the minimum achievable depth. Indeed, for larger values of $n$ we can sharpen the bound by setting $s=3$, which gives $D(n)\leq 3\lceil\log_{3}(n)\rceil-2\in\frac{3}{\log(3)}\log(n)+O(1)\approx 1.8928\log(n)+O(1)$. For any $n$, the actual minimum depth can be computed with dynamic programming: $\min D(n)=\min_{2\leq s\leq n/2}\{s+\min D(\lceil\frac{n}{s}\rceil-1)\}$.
@@ IV Quantum adders
Here we mention how to apply the proposed prefix circuits to quantum adders. Such circuits have received significant attention over the last three decades since they serve as building blocks for important problems such as discrete logarithm and integer factoring [24, 1, 31, 22, 30, 29, 7, 32]. Here we follow the methodology of [7, 32], where the main idea is to use a prefix circuit for carry propagation/generation (carry look-ahead framework). Circuits are evaluated on the following metrics:
@@
The maximum number of Toffoli gates in a path from an input to an output.
@@
The goal is to minimize the Toffoli depth/count by carefully overlapping different independent Toffoli layers, while still using as few auxiliary qubits as possible. Figure 3 shows three cases where Toffoli gates are used in the quantum circuit for propagation and generation. We refer also to [32, Sections 3 and 4] for a broad overview of the techniques and details of the specific operations.
@@ 3. $k:\bm{g}(i)\leftarrow\mathcal{T}\Big(\bm{g}(i-1),\bm{p}_{t-1}(i),\bm{g}(i)\Big)$ $i\in[s-1:s:n-1]$
The total number of Toffoli gates is bounded as follows. In the first two propagation layers, $\bm{p}_{0}$ and $\bm{p}_{1}$, there are a total of $3n/2$ gates, and additional $n/2$ gates for the un-computation of $\bm{p}_{1}$. In each recursive step $t=1,\ldots,\lceil\log_{s}(n)\rceil-1$, there are at most $n/s^{t}$ gates for $\bm{g}^{(t+1)}$, at most $n/s^{t+1}$ gates for $\bm{p}^{(t+2)}$, at most $n/2^{t}$ gates for $\bm{g}^{(t)}$, and finally at most $n/2^{t+1}$ gates for the uncomputation of the $\bm{p}^{(t+2)}$. This gives a total of at most $3n/2^{t}$ gates at level $t$. Summing for all $t=1,\ldots,\lceil\log(n)\rceil-1$ gives at most $3n$ Toffoli gates. ∎
Acknowledgements
We thank Igor Sergeev for helpful discussions and for indicating important references.